# Detect, Investigate, Judge and Determine: A Novel LLM-based Framework for Few-shot Fake News Detection

### v1

@@ Appendix A Prompts
In this section, we illustrate the prompts utilized in our DAFND methodology, which can serve as a valuable resource for future research in this area.
A.1 Detection Prompt
As a news keyword extractor, your task is to extract the $N$ most important keywords from a given news text. The keywords should include when, where, who, what, how and why the news happened. Please give me the six keywords only. My first suggestion request is {Target News Document}.
A.2 Inside Judge Prompt
@@
I need your assistance in evaluating the authenticity of a news article. I will provide you the news article and additional information about this news. You have to answer that [This is fake news] or [This is real news] in the first sentence of your output and give your explanation about [target news].
I will give you some examples of news. Your answer after [output] should be consistent with the following examples:
[example 1]:
[input news]: [news title: {}, news text: {}, news tweet: {}]
[output]: [This is {} news]
[target news]:
[input news]: [news title: {}, news text: {}, news tweet: {}]
@@
A.3 Outside Judge Prompt
I need your assistance in evaluating the authenticity of a news article. I will provide you the news article and additional information about this news. Please analyze the following news and give your decision. The first sentence of your [Decision] must be [This is fake news] or [This is real news].
The news article is: {}.
The additional information is: {}.
[Decision]:
A.4 Determination Prompt
@@
The news article is: news title: {}, news text: {}, news tweet: {}.
@@
Some people believe that {}, their explanation is: {}.
Others believe that {}, their explanation is: {}.
Please judge their opinion and give your decision. The first sentence after [Explanation] must be [This is fake news] or [This is real news], and then give your explanation.
@@
Appendix B Bad Case Analysis
In this section, we illustrate the bad cases that DAFND struggles with, with a goal to analyze its shortcomings and possible improvement directions.
Figure 5: The bad case of DAFND on the Gossipcop dataset ($K$=100).
As illustrated in Figure 5, a particular failure occurs when there are no relevant samples in the few-shot training set for retrieval. Consequently, the Inside Investigation and Inside Judge modules fail to work effectively. Meanwhile, although Outside Investigation retrieves a proof supporting the given news, Outside Judge still classifies it into Fake category because it claims the necessity of further confirmation. This bad case reveals two potential directions for improving our model:
For one thing, as introduced in Section 1, DAFND retrieves similar demonstrations from the training set in response to the Understanding Ambiguity problem. This relies on the assumption that valuable samples can be found in the training set, which is not always the case. As a consequence, the failure in Figure 5 occurs. Therefore, how to discriminate and mitigate the impact of such circumstances is a crucial direction for improving the DAFND design.
For another, as discussed in the Limitation section, we only employ the LLM to conduct inference, which cannot fully exploit LLMs’ powerful capabilities. In Figure 5, although Outside Investigation retrieves valuable information, the prediction of Outside Judge still goes wrong. In fact, despite the great reasoning ability of these general LLMs, they are not competent in news-related domains and are not sufficiently familiar with the specific expression characteristics. Thence, we would like to adopt fine-tuning techniques to adapt LLMs for the news corpus, which we believe could bring a positive effect to DAFND.

### v2

@@ Appendix C Prompts
In this section, we illustrate the prompts utilized in our DKFND methodology, which can serve as a valuable resource for future research in this area.
C.1 Detection Prompt
There is a news article and a list of concepts related to it. Your task is to determine how relevant these keywords are to the given news. Assign each keyword a relevancy score between 0 and 1, where 1 indicates the keyword is highly relevant, and 0 indicates it is not relevant at all."
Instructions:
1. Analyze the content of the news article to understand its main topics and themes.
2. Compare each concept to the news content and assign a relevancy score between 0 and 1.
3. Output the concepts and their scores in reverse order (from the most relevant to the least relevant).
4. Don’t return your explanation about the concepts and their scores.
Output Format:
[[\"concept 1\", <score>], [\"concept 2\", <score>],..., [\"concept N\", <score>]]
Example Input:
News Article:
Scientists have discovered a new method to reduce carbon emissions using advanced nanotechnology. This breakthrough could significantly impact efforts to combat climate change globally.
concepts:
Global warming, Carbon emissions, Cryptocurrency, Nanotechnology, Climate change
Example Output:
[[\"Global warming\", 0.8], [\"Carbon emissions\", 1.0], [\"Cryptocurrency\", 0.0], [\"Nanotechnology\", 1.0], [\"Climate change\", 0.9]]
Input:
News Article:
{Target News Document}.
concepts:
{Target News Concepts}.
Your Output Format:
[[\"concept 1\", #score#], [\"concept 2\", #score#],..., [\"concept N\", #score#]]
Please replace #score# in Your Output Format with the relevancy score of corresponding word:
C.2 Inside Judge Prompt
I need your help determining the reliability of some passages, and I’ve provided some relevant entities mentioned in the text. Here are some hints:
1. **Entity Accuracy**: Check if the entities (e.g., names, attributes, roles) match the passages I’ve provided. Note that we do not require all relevant entities to appear in the passage, you only need to verify that the entities present in the passage do not conflict with the entities provided. You also do not need to check the relationships between the entities. However, if none of the entities appear, then the passage should be unreliable.
Please output your reason and index of the two most relevant passage in the following JSON format:
{"reason": <your explanation for the decision>,"index": [<index of passage which is the most relevant>, <index of passage which is the second relevant>]}
Here are the Passages:
{Retrieved News Documents}.
Here are the relevant entities:
{Current News Concepts}.
Output:
{"reason": #Explanation#,"index": [#Index 1#, #Index 2#]}
Please replace #Explanation# with your explanation for the decision and replace #Index 1#, #Index 2# with index of passage which is the first relevant and second relevant.
C.3 Outside Judge Prompt
## Task Name:
Extract Relationships between Provided Entities from a Text.
## Task Description:
Your task is to extract triples that represent relationships between a set of **predefined entities** based on the provided text. The subject and object of each triple must come **only** from the list of entities that I provide. If no meaningful relationships are found between the entities, return **None**.
### Task Input:
1. A piece of text.
2. A list of possible entities.
### Task Output:
A list of triples in JSON format, or **None** if no valid relationships are found.
### JSON Format for Triples:
[{"subject": "subject_name", "predicate": "predicate_relation", "object": "object_name"}, {"subject": "subject_name","predicate": "predicate_relation","object": "object_name"}]
### Guidelines:
- The **subject** and **object** of the triple must come from the list of predefined entities. Do not extract any other words as entities.
- If there are no valid relationships between the provided entities, return **None**.
### Examples:
#### Example 1:
**Input Text:**
"Paris is the capital of France."
**Entities Provided:**
["Paris", "France"]
**Output JSON:**
[{"subject": "Paris", "predicate": "is the capital of", "object": "France"}]
#### Example 2:
**Input Text:**
"Einstein developed the theory of relativity."
**Entities Provided:**
["Einstein", "the theory of relativity"]
**Output JSON:**
["subject": "Einstein", "predicate": "developed", "object": "the theory of relativity"]
#### Example 3:
**Input Text:**
"She is a member of the organization."
**Entities Provided:**
["the organization"]
**Output JSON:**
None
#### Example 4:
**Input Text:**
"The Nile is the longest river in the world, flowing through Egypt."
**Entities Provided:**
["The Nile", "Egypt"]
**Output JSON:**
[{"subject": "The Nile", "predicate": "flows through", "object": "Egypt"}]
### Real Data:
**Input Text:**
{Target News Document}.
**Entities Provided:**
{Target News Concepts}.
**Output JSON:**
[{"subject": ##predefined_entity##, "predicate": #Relation#, "object": ##predefined_entity##}, {"subject": ##predefined_entity##, "predicate": #Relation#, "object": ##predefined_entity##}, {"subject": ##predefined_entity##, "predicate": #Relation#, "object": ##predefined_entity##}]
Please replace ##predefined_entity## in **Output JSON** of Real Data with entities from Entities Provided and replace #Relation# in **Output JSON** of Real Data with meaningful relationship between entities.
I do not need you give me code, return triples in JSON format, or **None** if no valid relationships are found.
C.4 Inside Determination Prompt
You are an expert of news authenticity evaluation. As an expert of news authenticity evaluation, you should analyze and evaluate the authenticity of news.
I need your assistance in evaluating the authenticity of a news article. Please focus on whether the content of the article is true, not on the genre and rhetoric of the article.
I will provide you the news article and additional information about this news.
@@ C.4 Inside Determination Prompt
You have to answer that [This is fake news] or [This is real news] in the first sentence of your output and give your explanation about [target news].
[example i]:
[input news]: [{}]
[output]: [This is real/fake news]
[target news]: [{}]
[input news]: [{}]
@@
C.5 Outside Determination Prompt
I need your assistance in evaluating the authenticity of a news article. I will provide you the news article and additional information about this news. Please analyze the following news and give your decision and reason. The first sentence of your [Decision and Reason] must be [This is fake news] or [This is real news], and then give reason.
The news article is:
{Target News Document}.
The additional information is:
{Outside Information}.
[Decision and Reason]:
C.6 Integrated Determination Prompt
@@
Here are the news article:
{Target News Document}.
@@
Some people believe that {}, their explanation is: {}
Others believe that {}, their explanation is: {}
Please judge their opinion, evaluate the authenticity of the news article by analyzing its logic, reliability, and factual accuracy, and finally give your decision.
The first sentence after [Explanation] must be [This is fake news] or [This is real news], and then give your explanation.
@@
C.7 Simplified Integrated Determination Prompt for Ablation Study
Here are the news article:
{Target News Document}.
There are two different views on this news article.
Some people believe that {}, their explanation is: {}
Others believe that {}, their explanation is: {}
Please give your decision.
The first sentence after [Explanation] must be [This is fake news] or [This is real news], and then give your explanation.
[Explanation]: